# A system of NLS arising in optical material without Galilean symmetry

### v1

@@ 4 Global existence
$S_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)<\mu_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c},$
$N_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0.$
@@
$\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\in A^{+}_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c}.$
@@ Appendix A Appendix
$\int|\nabla\psi|^{2}+3\gamma\omega\int|\psi|^{2}+(i\gamma c\cdot\nabla\psi,\psi)-9\int|\psi|^{2}-2\int|\psi|^{2}|\phi|^{2}-\frac{1}{9}\int\phi^{3}\bar{\psi}=0.$
@@
$\int|\nabla\phi|^{2}+\omega\int|\phi|^{2}+(ic\cdot\nabla\phi,\phi)+\int|\nabla\psi|^{2}+3\gamma\omega\int|\psi|^{2}+(i\gamma c\cdot\nabla\psi,\psi)$
@@
$-\frac{N-2}{2}\int|\nabla\psi|^{2}-3\gamma\omega\frac{N}{2}\int|\psi|^{2}-\gamma\frac{N-1}{2}(\psi,ic\cdot\nabla\psi)$
@@
$-\frac{N-2}{2}\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)-\frac{N}{2}\int\left(\omega|\phi|^{2}+3\gamma\omega|\psi|^{2}\right)$
@@ Acknowledgments.
This work was supported by the National Natural Science Foundation of China (12301090 and 11801519) and Fundamental Research Program of Shanxi Province NO.202203021222126. Part of this work was done by Y.L. during his postdoctoral studies at Central China Normal University, and he would like to thank the professor Shuangjie Peng for fruitful discussions and constant encouragement.

### v2

@@ 4 Global existence
$S_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)<\mu_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c,0},$
$N_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0.$
@@
$\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\in A^{+}_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c,0}.$
@@ Appendix A Appendix
$\int|\nabla\psi|^{2}+(3\gamma\omega-3\gamma+\alpha)\int|\psi|^{2}+(i\gamma c\cdot\nabla\psi,\psi)-9\int|\psi|^{2}-2\int|\psi|^{2}|\phi|^{2}-\frac{1}{9}\int\phi^{3}\bar{\psi}=0.$
@@
$\int|\nabla\phi|^{2}+\omega\int|\phi|^{2}+(ic\cdot\nabla\phi,\phi)+\int|\nabla\psi|^{2}+(3\gamma\omega-3\gamma+\alpha)\int|\psi|^{2}+(i\gamma c\cdot\nabla\psi,\psi)$
@@
$-\frac{N-2}{2}\int|\nabla\psi|^{2}-(3\gamma\omega-3\gamma+\alpha)\frac{N}{2}\int|\psi|^{2}-\gamma\frac{N-1}{2}(\psi,ic\cdot\nabla\psi)$
@@
$-\frac{N-2}{2}\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)-\frac{N}{2}\int\left(\omega|\phi|^{2}+3\gamma\omega|\psi|^{2}\right)-\frac{N}{2}(\alpha-3\gamma)\int|\psi|^{2}$
@@ Acknowledgments.
This work was supported by the National Natural Science Foundation of China (12301090) and Fundamental Research Program of Shanxi Province NO.202203021222126. Part of this work was done by Y.L. during his postdoctoral stay at Central China Normal University, and he would like to thank the professor Shuangjie Peng for fruitful discussions and constant encouragement.